# ACPO: AI-Enabled Compiler-Driven Program Optimization

### v2

@@ 3.2. ACPO Inference
The compilation process involving ML inference is depicted in Figure 2b. The input program is supplied to Clang to be translated into LLVM Intermediate Representation (IR). The LLVM IR is then passed through a sequence of passes and eventually reaches a pass that uses ACPOModel for its profitability analysis. The pass proceeds normally to confirm the legality of the transformation; when it needs to decide the profitability of the transformation, the ML framework is engaged. The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction. This step sets up the interface to the underlying ML framework, such as TensorFlow (Abadi et al., 2016). Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework. The ML framework is then instructed to run inference on the model, and the inference result is returned to the compiler. If the inference result is such that the transformation should be applied to a given code region, the transform pass then continues and applies the transformation. Otherwise, the pass gracefully rejects the option to transform the code and proceeds to either the next code region to be analyzed or terminates. In some cases, the inference result can contain a numeric value, e.g. the predicted optimal unroll factor for a given loop. To accomplish the tasks of training and inference we used several key components, which we describe in greater detail in the following subsections.
3.3. Compiler
The compiler of choice is LLVM (Lattner, 2008) version 12, which at a high level comprises the C/C++ frontend Clang, the mid-end optimizer, and a target-specific backend to generate executable code. The compiler first translates a program into LLVM IR, and then proceeds to optimize the IR via a sequence of passes. The passes are applied in sequence and at various scopes (module, function, loop, basic block). Once a pass has finished transforming a code region, appropriate internal data structures, including IR, are updated and the pass either proceeds to the next code region or terminates to enable the next pass to proceed. Once the compilation is done, an executable is produced that can be run and measured. We generally refer to the sequence of passes as a pipeline, and we focus on keeping the pipeline unchanged. This is important because empirical evidence shows that the order of passes and composition of the compiler pipeline affect program performance (Ashouri et al., 2017). In this work we focus on improving decision making within a pass using ML, so changes to the pipeline ordering and optimization selection of passes are orthogonal and outside the scope of the work.
3.4. ACPOModel
The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference. The header file code snippet is left out of this version for brevity, but it contains the same functionality described in figures 1 and 2.
@@ 3.4. ACPOModel
This class is a simple abstraction of how the model is incorporated into the compiler flow. The class is instantiated and a reference to the ML Interface (ACPOMLInterface class) is provided to enable the class to communicate with the ML framework. This framework can be active in the background (as is the case in our implementation) or be created on an as-needed basis. In either case, the reference to the interface is provided to establish communication. The model is then configured to return a specific type of result as a part of the ACPOAdvice class. Then the features to be used as input for inference are specified as input to addFeature method. Finally, to perform inference the caller would invoke the getAdvice() method and act on the result provided.
The getAdvice() method takes several simple steps, most importantly, does prepare the input data by first invoking prepareModelInput() to ensure that model inputs are provided to the model correctly and then invoking the runModel() method. The runModel method establishes communication with the ML framework, loads the model if it is not yet loaded, sends input features in, and then requests that the inference be run. If at any point in time there is an error, the process is aborted and the runModel() method will return false, subsequently causing ACPOAdvice to set to nullptr. This will be a crucial check to verify that ML inference is completed and should be checked by the caller. It is worth noting that the prepareModelInput() and runModel() methods are virtual. The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
3.5. ML APIs & Framework
@@
An important part of the ACPOModel class is the reference to the ACPOMLInterface class. This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler. The code snippet is left out of this version for brevity, but it contains the same functionality described in figure 2.
The ACPOMLInterface class is virtual and it is meant to be implemented depending on how your particular ML framework needs are accessed. If it is a compiled library linked with your version of LLVM or a separate process (as was the case for this paper), this class abstracts the exchange of data between the compiler and the ML framework so that it is fundamentally irrelevant to how the ML framework actually works. To use this class, the ACPOModel invokes routines such as loadModel, freeModel, setCustomFeatures, and others to set up the model on the side of the ML framework. Then when the input data is ready, runModel is invoked to perform inference. The result is then retrieved by the ACPOModel via the getModelResult family of methods.
@@
In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e. os.mkfifo(). This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis. While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes. Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 1. Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects and in this work to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment.
4. ACPO Models
@@
In this section, we provide two use cases of leveraging ACPO framework on optimizing compiler passes. As we mentioned earlier, state-of-the-art approaches showcase the benefits of adapting ML models for coarse-grain and fine-grain tuning of a number of compiler optimization passes We speculate, ACPO can benefit any mid/back-end optimization/transformation pass with an internal decision process mechanism and a set of inputs/outputs. Historically, compiler developers have incrementally improved such optimization passes by means of algorithmic and heuristic approaches (Aho et al., 1986; Hall et al., 2009) and we believe leveraging ML to enhance the internal decision-making process of optimization passes merits an in-depth investigation and lays the groundwork for many future works. To this end, we choose Loop Unroll and Function Inlining as two of the most important compiler optimizations used in many state-of-the-art compilers, including LLVM.
4.1. ACPO Loop Unrolling Model
@@
Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018). There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018). A number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.
A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1. The textttgetAdviceML() method first retrieves a reference to the ML interface which is used to communicate with the ML framework. It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file. The method then sets the inputs to the model and invokes the runModel method to begin inference. The inference process takes place on the ML framework, and the runModel call is a blocking call that waits until the inference completes. When it does, a successful inference will return desired results, which can be retrieved using getModelResultI methods. However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed.
@@
4.2. ACPO Function Inlining
@@
In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining. Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers. In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018; Theodoridis et al., 2022). A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
We invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made. Also, we respected the high-priority overrides. We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 4. We instantiate an ACPOFIModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework, similar to what we showed earlier with loop unroll pass. Once the ACPOFIModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOFIModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into ”FI-ShouldInline” field, which is then obeyed by the function inlining pass and applied as directed. Note that in this work, similar to loop unroll pass, we only replace the profitability checks with ML models.
5. Experimental Evaluation
In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of cross-validation method on Polybench Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
ACPO Loop Unrolling Model
@@ 5. Experimental Evaluation
As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side. Additionally, a new ML model has been designed and deployed for this matter. The details of the design and testing are provided as follows. ACPOLUModel is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor. i.e., UP.Count values, given a loop. We have designed and implemented our model in PyTorch (Paszke et al., 2017), but once we train it, we convert the model into a TensorFlow (TF) (Abadi et al., 2016) frozen model and leverage TF’s inference mechanism to do inference at deployment. As mentioned earlier in Section 3.5, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match. The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed. The features are passed with std::vector of pairs to the ML side. Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR. The features used are shown in Table 3 and taken from our super class of ACPOFeatures which contains a large set of handcrafted features defined in different scopes, i.e., Module, Function, Call Graph, and Loops, so that users can easily add and register them into their optimization use cases. The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e. classes corresponding to unrolling count, is also depicted in Figure 3. The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label. We did this by using features with at least a 2% correlation.
Data Collection
We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are . The autotuner uses the list of loop unrolling count opportunities to determine optimal configurations by using different search techniques. Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file. This process was repeated for 100 iterations to generate our training data. Note that each iteration, depending on the number of loops an application has, generates multiple data points, hence the total number of training data is significantly higher than the number of iterations of tuning rounds. The data generation process takes around 2 days. The execution time of a number of Polybench applications can be too short to allow for a reliable measurement and to this end, we define a larger input dataset for those having an execution time of less than a second. This includes: gemver, jacobi-1d, bicg, atax, mvt, durbin, trisolv, gesummv, and deriche. In this work, we define the Loop Unrolling speedup as the following:
@@
where is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner. In this work, total program runtime and the are collected with Linux perf. All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core. All experiments are measured on an ARMv8.2-A (Kunpeng 920) architecture @ 2.6GHz on Linux. We run each benchmark three times and before each measurement run, we flush the system page cache to avoid any perturbance into the collection of our training data and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust. Additionally, we treat a single instance of iteration of data generation, normally having multiple loops and generating multiple samples, one set of runs meaning that we assign the Loop_speedup value to the whole set of rows generated as a result of the configuration. This way, we can filter out bad runs from good runs simply by pruning those causing a slowdown, or even having at least a certain amount of speedup values. This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
@@
We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table 1. As mentioned earlier under Section 4.1, Loop Unroll pass is registered twice at O3’s pipeline. We have implemented ACPO in a way that we load our ML models once per each module and keep the models alive in the lifetime of a module to avoid a repeated loading process of the models per each instance of inference. To this end, the other contributing factor is the number of modules per application and to a lesser extent the number of loops inside a module. The number of modules determines how many times the ML model needs to be loaded, which is an expensive task, and the number of loops determines the number of times Loop Unroll pass will be activated and ACPOLUModel will be deployed. ACPOLUModel is persistent and loaded once per module and we are working on making it even more optimized, i.e., loaded once per each application having multiple modules. For example. On Polybench’s bicg, we have two modules, namely, polybench.c and bicg.c. The former has a single loop and the latter, has 8, thus, the overall overhead would be 0.857 seconds.
@@
The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test. This ensures no data leakage occurs. Measured Top-1 accuracy of the ACPOLUModel’s predictions exhibits a geometric mean of 80% accuracy with a standard deviation of 7% on Polybench. In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower. Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques (Wei and Zou, 2019). We leave further investigation to our future works as they are outside the scope of this paper. Subsequent to establishing the high top-1 accuracy on a cross validation mode, we evaluate the performance of the ACPOLUModel by means of deploying it on Polybench and measuring the actual benefits of our ML model. Table 2 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012) on Aarch64. We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall. Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
@@
Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts. ACPO’s prediction provides for a full unroll of a couple of loops including a multiple runtime unrolling of hot loops in the kernel which matches the optimal result our autotuner found after an exhaustive search. O3 on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops. and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
5.1. ACPO Function Inlining Model
Similar to Loop Unrolling, Function inlining pass has been optimized by means of adding the necessary code to connect the interface with the ML side but instead of rebuilding an ML model from scratch, we decided to leverage an existing model to showcase the ease of use and the integration capabilities of ACPO when paired with an existing ML model. For this task, we use MLGOPerf (Ashouri et al., 2022) model.
5.1.1. Function Inlining Model
MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm. The model receives as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite is to be inlined. Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology. The model has extended MLGO (Trofin et al., 2021)’s code size reduction to optimize performance instead. As shown in Listing 4, getACPOAdvice() initiates the call to collect necessary features and finally returns the Boolean value of the decision of whether or not to inline a callsite.
5.1.2. Function Inlining Generalization Results
Table 3 showcases the experimental results of the deployed ACPOFIModel when tested with CBench. As expected, the results exhibit similar behavior to MLGOPerf’s experimental results since both are leveraging the same model, albeit differently. For instance, we have gained around 27% on security_sha and inspecting the decisions made by ACPOFIModel, we notice that LLVM’s O3 did not inline the two hot callsites namely, sha_update() and sha_final() whereas in ACPO, those two are inlined. A geometric mean showcases around 2% increase in performance by incurring around 15% code size bloat on Cbench.
5.2. Combined Results
Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench. An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).

### v3

@@ III-B ACPO Inference
The compilation process involving ML inference is depicted in Figure 2b. The input program is supplied to Clang to be translated into LLVM Intermediate Representation (IR). The LLVM IR is then passed through a sequence of passes and eventually reaches a pass that uses ACPOModel for its profitability analysis. The pass proceeds normally to confirm the legality of the transformation; when it needs to decide the profitability of the transformation, the ML framework is engaged. The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction. This step sets up the interface to the underlying ML framework, such as TensorFlow [49]. Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework. The ML framework is then instructed to run inference on the model, and the inference result is returned to the compiler. If the inference result is such that the transformation should be applied to a given code region, the transform pass then continues and applies the transformation. Otherwise, the pass gracefully rejects the option to transform the code and proceeds to either the next code region to be analyzed or terminates. In some cases, the inference result can contain a numeric value, e.g. the predicted optimal unroll factor for a given loop.
III-C ACPOModel
@@
The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to perform inference. The header file code snippet is left out of this version for brevity, but it contains the same functionality described in Figure 2. This class is a simple abstraction of how the model is incorporated into the compiler flow. The class is instantiated, and a reference to the ML Interface (ACPOMLInterface class) is provided to enable the class to communicate with the ML framework. The model is then configured to return a specific type of result as a part of the ACPOAdvice class. Then, the features to be used as input for inference are specified as input to addFeature method. Finally, to perform inference, the caller would invoke the getAdvice() method and act on the result provided. It is worth noting that the prepareModelInput() and runModel() methods are virtual to enable derivation of this class to represent a new custom models.
III-D ML APIs & Framework
The snippet for the ACPOMLInterface header is shown under the Appendix section in Listing 6.
@@ III-D ML APIs & Framework
An important part of the ACPOModel class is the reference to the ACPOMLInterface class. This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class and the compiler. The code snippet is left out of this version for brevity, but it contains the same functionality described in figure 2.
Communication
@@
In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e. os.mkfifo(). This approach allows us to completely separate the compiler from the ML framework, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis. While at the moment, the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes.
Feature Collection Class
We have implemented a ACPOCollectFeature class having around 130 handcrafted features in total at different scopes of LLVM’s IR, from which 80 are Function level, 35s are Loops, and around 15 are Module features. Users can easily select features individually or by their scope when creating ACPOModel classes.
IV ACPO Models
@@
In this section, we provide two use cases of leveraging ACPO framework on optimizing compiler passes. As we mentioned earlier, state-of-the-art approaches showcase the benefits of adapting ML models for coarse-grain and fine-grain tuning of a number of compiler optimization passes We speculate, ACPO can benefit any mid/back-end optimization/transformation pass with an internal decision process mechanism and a set of inputs/outputs. Historically, compiler developers have incrementally improved such optimization passes by means of algorithmic and heuristic approaches [50, 2] and we believe leveraging ML to enhance the internal decision-making process of optimization passes merits an in-depth investigation and lays the groundwork for many future works. To this end, we choose Loop Unroll and Function Inlining as two of the most important compiler optimizations used in many state-of-the-art compilers, including LLVM.
IV-A ACPO Loop Unrolling Model
@@
Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity [51, 30]. There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count [51, 52, 53, 54].
The snippet for the ACPOLUModel is shown under the Appendix section in Listing 7 A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 7. To leverage ML in unrolling, we invoked the ACPOLUModel after the loop unrolling legality checks had been made. Also, we respected the high-priority overrides. The highest level is set by the explicit use of unroll-count by the user, followed by the usage of #pragma clang loop unroll_count(N). We invoke ACPOLUModel object subsequent to these two priorities as shown in Listing 2.
@@
IV-B ACPO Function Inlining
@@
In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining. Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers. In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction [15, 55] and performance [8] and it was shown the optimal function inlining problem has an exponential increase in optimization space [30, 55]. A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 8 .
The snippet for the ACPOLFIModel is shown under the Appendix section in Listing 8
We invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made. Also, we respected the high-priority overrides. We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 3.
V Experimental Evaluation
Note that in an ML-guided (MLGO) approach such as ACPO, optimizing a pass overrides their existing profitability checks and thus, provides a unique opportunity to run the pass in the same predefined order the sequence of passes is run at LLVM’s O3. Thus, all evaluations here are based on comparing the performance of O3 with the vanilla or the ML-enabled passes. For the same reason and except for the performance comparison with MLGO work [15] under Table V, we are not evaluating ACPO with the many other aforementioned works as their method is outside the scope of an MLGO approach. To this end, we first evaluate our two passes individually, and finally, a combined evaluation is presented.
V-A ACPO Loop Unrolling Model
@@
We design our ACPOLUModel with a 5-layer Neural Net using ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor. i.e., UP.Count values, given a loop fed as input. The input to the model is a set of 30 handcrafted features selected from the ACPOCollectFeature class at the scope of Loops. The features are passed with std::vector of pairs to the ML side. Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc. However, we leverage handcrafted features derived from LLVM IR in this work. The Pearson [56] correlation analysis between the features and the target label, i.e. classes corresponding to unrolling count, is also depicted in Figure 3.
| No | Static Features | No | Static Features |
| 1 | PartialOptSizeThreshold | 16 | NumLoadInstPerLoopNest |
| 2 | AllowRemainder | 17 | NumStoreInstPerLoopNest |
| 3 | UnrollRemainder | 18 | TotLoopNestInstCount |
| 4 | AllowExpensiveTripCount | 19 | AvgNumLoadInstPerLoopNest |
| 5 | Force | 20 | AvgNumStoreInstPerLoopNest |
| 6 | TripCount | 21 | NumLoadInstPerLoop |
| 7 | MaxTripCount | 22 | NumStoreInstPerLoop |
| 8 | Size | 23 | TotLoopInstCount |
| 9 | InitialIVValueInt | 24 | AvgNumLoadInstPerLoop |
| 10 | FinalIVValueInt | 25 | AvgNumStoreInstPerLoop |
| 11 | StepValueInt | 26 | IsInnerMostLoop |
| 12 | NumPartitions | 27 | IsOuterMostLoop |
| 13 | IndVarSetSize | 28 | MaxLoopHeight |
| 14 | AvgStoreSetSize | 29 | TotBlocksPerLoop |
| 15 | AvgNumInsts | 30 | IsFixedTripCount |
Data Collection & Setup
@@
We used an autotuner to collect a total of 100000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are . In this work, we define the Loop Unrolling speedup as the fraction between the base runtime and the loop runtime. The total program runtime and the are collected with Linux perf. All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core. All experiments are measured on an ARMv8.2-A architecture @ 2.6GHz on Linux. We run each benchmark three times, and before each measurement run, we flush the system page cache to avoid any perturbance in the collection of our training data, and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust.
@@
We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table II. As mentioned earlier under Section IV-A, Loop Unroll pass is registered twice at O3’s pipeline. We have implemented ACPO in a way that we load our ML models once per each module and keep the models alive in the lifetime of a module to avoid a repeated loading process of the models per each instance of inference. To this end, the other contributing factor is the number of modules per application and to a lesser extent the number of loops inside a module. The number of modules determines how many times the ML model needs to be loaded, which is an expensive task, and the number of loops determines the number of times Loop Unroll pass will be activated and ACPOLUModel will be deployed. ACPOLUModel is persistent and loaded once per module and we are working on making it even more optimized, i.e., loaded once per each application having multiple modules. For example. On Polybench’s bicg, we have two modules, namely, polybench.c and bicg.c. The former has a single loop and the latter, has 8, thus, the overall overhead would be 0.857 seconds.
@@
The training is carried out using leave-one-out cross-validation, for which we exclude one application from the training data and use it later as the test, and this ensures no data leakage occurs. Measured Top-1 accuracy of the ACPOLUModel’s predictions exhibits a geometric mean of 80% accuracy with a standard deviation of 7% on Polybench. In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower. Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques [57]. We leave further investigation to our future works as they are outside the scope of this paper.
| Benchmark | Top-1 Accuracy | Benchmark | Top-1 Accuracy |
| 2mm | 88.411 % | gesummv | 80 % |
| 3mm | 84.254 % | gramschmidt | 75.466 % |
| adi | 76.543 % | lu | 85.971 % |
| atax | 57.5 % | ludcmp | 85.938 % |
| bicg | 63.295 % | mvt | 75.326 % |
| cholesky | 86.414 % | nussinov | 79.734 % |
| correlation | 84.091 % | seidel-2d | 80.909 % |
| covariance | 86.42 % | symm | 79.797 % |
| deriche | 82.481 % | syr2k | 84.858 % |
| doitgen | 75.487 % | syrk | 82.84 % |
| floyd-warshall | 82.864 % | trisolv | 69.225 % |
| gemver | 86.547 % | trmm | 90.333 % |
| Geomean: 80.196% | |||
@@
Table VI shows the ACPOLUModel’s results when deployed in a cross-validation mode. We observe that, on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall. Further investigation reveals that the Floyd-Warshall application has a seven-loop module, which makes it for a set of 23 tunable loop-nest regions. ACPO’s prediction provides for a full unroll of a couple of loops, including a multiple runtime unrolling of hot loops in the kernel. LLVM’s O3, on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops. Thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
Generalization
To further evaluate the generalizability of the model, we have tested ACPOLUModel on a set of heavily loop-oriented benchmarks. Table IV demonstrates our result for which we gain a 5.4% speedup over CoreMark, an average 3% including an 18% speedup over AMG (from Coral-2), and a 1.3% speedup on Graph500’s SSSP. Analyzing the results, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops. For instance, Coral-2’s AMG runtime unrolling of a hot loop located at hypre_CSRMatrix Module led to an increase in performance by almost 15%, which O3 failed to identify as profitable to unroll.
| Benchmarks | O3 | ACPO | SpUp | ||
| Time | Size | Time | Size | ||
| CoreMark | 20.783 | 74328 | 19.723 | 74328 | 1.054 |
| CLOMP | 142.267 | 78952 | 140.835 | 78952 | 1.010 |
| Quicksilver | 11.647 | 256072 | 11.657 | 527328 | 0.999 |
| STRIDE cac. | 19.707 | 72112 | 18.83 | 72112 | 1.047 |
| STRIDE str. | 44.842 | 71928 | 44.792 | 71928 | 1.001 |
| STREAM | 18.899 | 72992 | 18.786 | 72992 | 1.006 |
| AMG | 3.2922 | 564192 | 2.785 | 1292568 | 1.182 |
| LAMMPS | 2.92 | 3802424 | 2.826 | 7082568 | 1.033 |
| BFS | 7.275 | 539264 | 7.33 | 539240 | 0.992 |
| SSSP | 15.508 | 539832 | 15.308 | 605344 | 1.013 |
V-B ACPO Function Inlining Model
Instead of designing a novel ML model from scratch, we decided to leverage an existing model to showcase the ease of use and integration capabilities of ACPO when paired with an existing ML model. For this task, we use the MLGOPerf [8] model. MLGOPerf is a reinforcement learning model, and the model receives, as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite should be inlined. Table V showcases the experimental results of using ACPO framework for the deployment of ACPOFIModel, aka MLGOPerf’s, when tested with CBench. As expected, the results exhibit similar behavior to MLGOPerf’s experimental results since both leverage the same model.
V-C Combined Results
Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO [15].
| Benchmark | ACPO-FI | wrt. MLGO | ACPO-LU | Combined | ||||
| Sz% | Sp | Sz% | Sp | Sz% | Sp | Sz% | Sp | |
| auto_bitcount | 1 | 0.999 | 1 | 1.005 | 1.001 | 1.003 | 1 | 1.001 |
| auto_qsort1 | 1 | 0.994 | 1 | 1.003 | 1 | 0.982 | 1.91 | 0.989 |
| auto_susan_c | 1 | 0.999 | 1.567 | 1.122 | 3.685 | 1.035 | 19.797 | 1.066 |
| auto_susan_e | 1 | 1.023 | 1.567 | 1.04 | 3.685 | 1.019 | 19.797 | 1.026 |
| auto_susan_s | 1 | 0.98 | 1.567 | 1.008 | 3.685 | 1.038 | 19.797 | 0.995 |
| bzip2d | 1.441 | 0.991 | 1.239 | 1.021 | 2.224 | 1.01 | 3.541 | 1.014 |
| bzip2e | 1.441 | 1.007 | 1.189 | 1.041 | 2.35 | 1.001 | 3.546 | 1.015 |
| cons_jpeg_c | 1.288 | 1.017 | 1.2 | 1.005 | 5.385 | 0.986 | 5.086 | 1.027 |
| cons_jpeg_d | 1.289 | 0.923 | 1.19 | 1.335 | 5.4 | 1.073 | 4.807 | 1.026 |
| cons_lame | 1.437 | 1.027 | 1.137 | 1.007 | 4.501 | 1.03 | 4.729 | 1.034 |
| cons_mad | 1.217 | 1.071 | 1.16 | 1 | 4.321 | 0.99 | 3.193 | 0.968 |
| cons_tiff2bw | 1.18 | 1.002 | 1.351 | 1.23 | 4.502 | 0.943 | 2.471 | 0.987 |
| cons_tiffmedian | 1.18 | 1.017 | 1.048 | 1.056 | 1 | 0.97 | 3.021 | 0.983 |
| net_dijkstra | 1 | 0.993 | 1 | 0.996 | 1 | 0.99 | 1 | 1.037 |
| net_patricia | 1 | 0.98 | 1.052 | 0.99 | 1 | 0.974 | 1 | 1.004 |
| off_ghostscript | 1.275 | 1 | 1.16 | 1.04 | 1 | 1 | 1.12 | 1.001 |
| office_ispell | 0.996 | 1 | 1.17 | 1.02 | 1 | 1 | 1 | 1 |
| office_rsynth | 1 | 1 | 1.12 | 1.04 | 2.024 | 1.01 | 2.024 | 1.011 |
| office_strings | 1 | 1.006 | 1 | 1 | 1 | 1 | 1 | 1 |
| sec_blowfish_d | 1 | 1.002 | 1 | 1.002 | 1 | 0.978 | 1 | 0.965 |
| sec_blowfish_e | 1 | 1.003 | 1.39 | 1.029 | 1 | 1.015 | 1 | 0.981 |
| security_pgp_d | 1.861 | 0.993 | 1.12 | 1 | 1 | 0.991 | 13.144 | 1 |
| security_pgp_e | 1.861 | 1.017 | 1.12 | 1.042 | 1 | 1.006 | 13.144 | 1.006 |
| sec_rijndael_d | 1 | 0.99 | 1.39 | 1.039 | 3.823 | 0.966 | 1 | 0.966 |
| sec_rijndael_e | 1 | 1.001 | 1.09 | 1.035 | 3.823 | 0.974 | 1 | 0.995 |
| security_sha | 0.999 | 1.27 | 1.009 | 1.132 | 1 | 1.174 | 0.999 | 1.501 |
| tele_adpcm_c | 1 | 1 | 0.996 | 1.001 | 1 | 1.008 | 1 | 1.038 |
| tele_adpcm_d | 1 | 0.997 | 1 | 1 | 1 | 1.132 | 1 | 1.138 |
| tele_CRC32 | 1 | 1.094 | 1 | 1.066 | 1 | 1.042 | 1 | 1.112 |
| telecom_gsm | 1.81 | 1.076 | 1.05 | 1.04 | 1.815 | 1.15 | 1.81 | 1.117 |
| Geomean | 1.152 | 1.021 | 1.161 | 1.041 | 1.818 | 1.01 | 2.442 | 1.024 |